# A Sanity Check for AI-generated Image Detection

### v1

@@ 3.2.3 Dataset Annotation
At this stage, we establish an annotation platform and recruit 20 human workers to manually label each of the AI-generated images for their category and realism. For categorization, annotators are instructed to assign each image to one of four major categories: human, animal, object, and scene. Regarding realism assessment, workers are tasked with labeling the images as either Real or AI-generated, based on the criterion of “whether this image could be taken with a camera”. It’s important to note that as the annotators are not informed whether the images are generated by AI algorithms beforehand, those have been misclassified as real can thus be considered to pass the ‘Perception Turing Test’ and labeled as “highly realistic”. Subsequently, we retain only those images judged as “highly realistic”. Similarly, for real images, we follow the same procedure, retaining only those belonging to the four predefined categories, as we have done for AI-generated images.
4 Architecture
@@ 4.1 Patchwise Feature Extraction
To acquire the highest and lowest image patches, we use the complexity of the frequency components as an indicator. From this, we design a simple yet effective scoring mechanism. Specifically, firstly, we design $K$ different band-pass filters:
@@
where $c$ is the number of patch channels. Thus, we acquire the grades of all patches $G=\{G^{1},G^{2},...,G^{n}\}$, and then we sort them to identify the highest and lowest frequency patches.
@@
Patchwise Feature Encoder. Next, firstly, these patches are resized to a size of $256\times 256$. Secondly, they are input into the SRM [21] to extract their noise pattern. Lastly, these features are input into two ResNet50 [25] network ($f_{1}(\cdot)$ and $f_{2}(\cdot)$) to obtain the final feature map $F_{\text{max}}=\{f_{1}(X_{\text{max}_{1}}),f_{1}(X_{\text{max}_{2}}),...,f_{1}(X_{\text{max}_{k}})\}$, $F_{\text{min}}=\{f_{2}(X_{\text{min}_{1}}),f_{2}(X_{\text{min}_{2}}),...,f_{2}(X_{\text{min}_{k}})\}$. We acquire the highest frequency embedding and lowest frequency embedding on the mean-pooled feature:
@@ 4.2 Semantic Feature Embedding
To capture the rich semantic features within images, such as object co-occurrence and contextual relationships, we compute the visual embedding for input image with an off-the-shelf visual-language foundation model. Specifically, we adopt the ConvNeXt-based OpenCLIP model [31] to get the final feature map ($v\in\mathbb{R}^{h\times w\times c}$). To capture the global contexts, we append a linear projection layer followed by mean spatial pooling, $F_{\text{s}}=\text{avgpool}(g(v))$.
@@ 4.3 Discriminator
To distinguish between AI-generated images and real images, we utilize a mixture-expert-model for the final discrimination. At low-level, we take the average of the highest frequency feature $F_{\text{\text{max}}}$ and the lowest frequency feature $F_{\text{\text{min}}}$. Then, we channel-wisely concatenate the representations between it and high-level embedding $F_{s}$. At last, the features are encoded into MLP to acquire the score, $y=f([\text{avgpool}(F_{\text{max}},F_{\text{min}});F_{\text{s}}])$, where $f(\cdot)$ denotes the MLP consisting of a linear layer, GELU [26] and classifier, [;][;][ ; ] refers to the operation of channel-wise concatenation.
@@ 5 Experiments
Table 2: Benchmark1. Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). DIRE-D indicates this result comes from DIRE detector trained over fake images generated by ADM following its official setup [63]. DIRE-G indicates this baseline is trained on the same ProGAN training data as others. GAN-Average and DM-Average are averaged over the first 8 and the last 8 test sets, respectively. The best result and the second-best result are marked in bold and underline, respectively.
Generator CNNSpot [62] FreDect [20] Fusing [32] GramNet [40] LNP [37] LGrad [59] UnivFD [49] DIRE-G [63] DIRE-D [63] PatchCraft [70] Ours
ProGAN 100.00 99.36 100.00 99.99 99.67 99.83 99.81 95.19 52.75 100.00 99.99
StyleGAN 90.17 78.02 85.20 87.05 91.75 91.08 84.93 83.03 51.31 92.77 99.64
BigGAN 71.17 81.97 77.40 67.33 77.75 85.62 95.08 70.12 49.70 95.80 83.95
CycleGAN 87.62 78.77 87.00 86.07 84.10 86.94 98.33 74.19 49.58 70.17 98.48
StarGAN 94.60 94.62 97.00 95.05 99.92 99.27 95.75 95.47 46.72 99.97 99.91
GauGAN 81.42 80.57 77.00 69.35 75.39 78.46 99.47 67.79 51.23 71.58 73.25
StyleGAN2 86.91 66.19 83.30 87.28 94.64 85.32 74.96 75.31 51.72 89.55 98.00
WFIR 91.65 50.75 66.80 86.80 70.85 55.70 86.90 58.05 53.30 85.80 94.20
ADM 60.39 63.42 49.00 58.61 84.73 67.15 66.87 75.78 98.25 82.17 93.45
Glide 58.07 54.13 57.20 54.50 80.52 66.11 62.46 71.75 92.42 83.79 95.09
Midjourney 51.39 45.87 52.20 50.02 65.55 65.35 56.13 58.01 89.45 90.12 77.20
SD v1.4 50.57 38.79 51.00 51.70 85.55 63.02 63.66 49.74 91.24 95.38 93.00
SD v1.5 50.53 39.21 51.40 52.16 85.67 63.67 63.49 49.83 91.63 95.30 92.85
VQDM 56.46 77.80 55.10 52.86 74.46 72.99 85.31 53.68 91.90 88.91 95.16
Wukong 51.03 40.30 51.70 50.76 82.06 59.55 70.93 54.46 90.90 91.07 93.55
DALLE2 50.45 34.70 52.80 49.25 88.75 65.45 50.75 66.48 92.45 96.60 96.60
GAN-Average 87.94 78.78 84.21 84.87 86.76 85.28 77.39 50.79 91.90 88.21 93.43
DM-Average 53.61 49.28 52.55 52.48 80.91 65.41 59.97 92.28 64.95 90.42 92.11
Average 70.78 64.03 68.38 68.67 83.84 75.34 78.43 68.68 71.53 89.31 92.77
Table 3: Benchmark2. Accuracy (%) of different baselines (columns) in detecting real and fake images from different generators (rows). These methods are trained on real images from ImageNet and fake images generated by SD v1.4 and evaluated over eight generators. The best result and the second-best result are marked in bold and underline, respectively.
Testing Dataset Average
Midjourney SD v1.4 SD v1.5 ADM GLIDE Wukong VQDM BigGAN
ResNet-50 [25] 54.90 99.90 99.70 53.50 61.90 98.20 56.60 52.00 72.09
DeiT-S [60] 55.60 99.90 99.80 49.80 58.10 98.90 56.90 53.50 71.56
Swin-T [41] 62.10 99.90 99.80 49.80 67.60 99.10 62.30 57.60 74.78
CNNSpot [62] 52.80 96.30 95.90 50.10 39.80 78.60 53.40 46.80 64.21
Spec [68] 52.00 99.40 99.20 49.70 49.80 94.80 55.60 49.80 68.79
F3Net [51] 50.10 99.90 99.90 49.90 50.00 99.90 49.90 49.90 68.69
GramNet [40] 54.20 99.20 99.10 50.30 54.60 98.90 50.80 51.70 69.85
DIRE [63] 60.20 99.90 99.80 50.90 55.00 99.20 50.10 50.20 70.66
UnivFD [49] 73.20 84.20 84.00 55.20 76.90 75.60 56.90 80.30 73.29
GenDet [73] 89.60 96.10 96.10 58.00 78.40 92.80 66.50 75.00 81.56
PatchCraft [70] 79.00 89.50 89.30 77.30 78.40 89.30 83.70 72.40 82.30
Ours 79.38 99.74 99.76 78.54 91.82 98.65 80.26 66.89 86.88
5.1 Experimental Setup
@@
Detectors. We evaluate 9 off-the-shelf detectors including CNNSpot [62], FreDect [20], Fusing [32], GramNet [40], LNP [37], LGrad [59], UnivFD [49], DIRE [63] and PatchCraft [70] for comparsion. More details can be found in Appendix.
@@
Metrics. In accordance with existing AI-generated detection arpproaches [62, 61, 71], we report both classification accuracy (Acc) and average precision (AP) in our experiments. All results are averaged over both real and AI-generated images unless otherwise specified. We primarily report Acc for evaluation and comparison in the main paper, and AP results are presented in the Appendix.
5.2 Benchmarks
To comprehensively evaluate the generalization ability of existing approaches, we detail three benchmarks, (i) AIGCDetectBenchmark [62], (ii) GenImage [74], and (iii) our Chameleon dataset, where (i) and (ii) belong to Setting-I and (iii) belongs to Setting-II as summarized in Sec. 3.1. More detailed statistics of these benchmarks are shown in Appendix.
Benchmark 1 ($\mathcal{B}_{1}$): We follow the widely-used AIGCDetectBenchmark [62, 74, 70], which aims to constrain the detector to learn a universal fingerprint from one generator and generalize to others. Specifically, the training dataset consists of 360k real images from LSUN [67] and 360k fake images generated only by ProGAN [33] with the same classes. In evaluation, we follow PatchCraft’s [70] setup, which covers most existing foundational generators, collected from ForenSynths [62] and GenImage [74]. The test dataset includes 16 different generative models, with both GAN-based and DM-based architectures, providing a comprehensive evaluation of detection performance on these mainstream generators commonly utilized in both research and industry. This includes advanced GAN-based generators like StyleGAN2 [35] and commercial DM-based APIs like Midjourney [4].
Benchmark 2 ($\mathcal{B}_{2}$): GenImage is a newly proposed benchmark and has been evaluated in recent studies [70, 14, 23]. The two major improvements over $\mathcal{B}_{1}$ include incorporating advanced DM-based generators and expanding the training classes from LSUN’s 20 classes to ImageNet’s 1,000 classes. In particular, GenImage includes ImageNet as real images and generates images with ImageNet’s labels as conditions. These images are generated using 8 generators, including BigGAN [13], ADM [17], Glide [46], Midjourney [4], SD v1.4 [6], SD v1.5 [6], VQDM [24], and Wukong [10]. Each generator corresponds to a training and testing set consisting of an equal number of real and generated images, resulting in 8 training sets and 8 testing sets. Following GenImage’s setup, we utilize SD v1.4 as the training dataset and report the averaged results over 8 test datasets.
Benchmark 3 ($\mathcal{B}_{3}$): To test the models on our Chameleon testset, we train them under different training setups: (i) we adopt the training data from $\mathcal{B}_{1}$ as it is widely used among these baselines, which has demonstrated excellent generalization ability across on the test of $\mathcal{B}_{1}$; (ii) the models are trained on the training data from $\mathcal{B}_{2}$; (iii) the models are trained on the generated images from 8 generators in GenImage, including both the classic GAN-based models (e.g., BigGAN) and the latest DM-series models (e.g., SD and Midjourney).
5.3 Comparison with State-of-the-arts
@@
On Benchmark 1: In Table 2, the quantitative results showcase the classification accuracies across various methods and generators within $\mathcal{B}_{1}$. In our evaluation, all methods except DIRE-D were trained exclusively on ProGAN’s generations.
@@
Our approach demonstrates a notable improvement over the existing SOTA method, PatchCraft, achieving a 3.5% higher average accuracy. Among the baseline methods, UnivFD incorporates CLIP semantic features for AI-generated image detection. It exhibits effectiveness in detecting GAN-generated images but experiences a performance drop when applied to DM-generated images, showing that with the growing generation quality, images generated by diffusion model tend not to lead discernible artifacts in semantics, as illustrated in Fig. 1 (a). While our method integrates semantic, low-frequency and high-frequency information at the feature level, leading to improved detection performance for both GAN-based (5.2% increase) and DM-based (1.7% increase) images compared to the SOTA method.
@@
On Benchmark 2: While conducting experiments on $\mathcal{B}_{2}$, where all models were trained on SD v1.4 and evaluated across 8 up-to-date generators. Table 3 summarizes the results, highlighting our method’s superior performance over the current state-of-the-art method, PatchCraft, by 4.6% in terms of average accuracy. It is worth noting the architectural similarities between SD v1.5, Wukong, and SD v1.4, as highlighted by GenImage [74]. This resemblance enables models to achieve nearly perfect accuracy, approaching 100% on such test sets. Consequently, the assessment of generalization performance on other generators becomes crucial, for example, Midjourney, ADM, and Glide, where our model demonstrates either the best or second-best results, with an average accuracy of 86.88%.
@@
Table 4: Benchmark3. Accuracy (%) of different detectors (rows) in detecting real and fake images of Chameleon testset (rows). For each training dataset, the first row indicates the average Acc evaluated on the Chameleon testset, and the second row gives "fake image Acc / real image Acc" for detailed analysis.
@@
Training Dataset CNNSpot [62] FreDect [20] Fusing [32] GramNet [40] LNP [37] LGrad [59] UnivFD [49] DIRE [63] PatchCraft [70] Ours
ProGAN 49.88 50.42 49.90 52.21 50.03 50.63 50.51 51.37 47.30 49.55
0.08/99.67 13.72/87.12 0.01/99.79 4.76/99.66 0.09/99.97 1.70/99.55 3.18/97.83 3.25/99.48 1.78/92.82 0.63/98.46
SD v1.4 56.25 49.97 49.98 55.58 48.79 55.49 58.03 53.77 49.71 61.54
13.86/98.63 1.37/98.57 0.00/99.96 17.65/93.50 0.57/97.01 11.44/99.54 74.97/41.09 11.86/95.67 3.07/96.35 40.61/82.46
All GenImage 54.56 50.22 50.00 53.41 52.21 52.48 63.54 50.91 48.96 58.73
9.86/99.25 0.89/99.55 0.02/99.98 8.23/98.58 7.72/96.70 5.19/99.77 85.52/41.56 2.09/99.73 1.39/96.52 22.40/95.06
On Benchmark 3: As emphasized in Sec. 1, we argue that success on the existing public benchmarks may not truly reflect the progress of AI-generated image detection, as the testsets are typically randomly sampled from generative models.
To mitigate potential biases introduced by training setups, such as generator types and image categories, we assess the performance of existing baselines in diverse training settings. Despite their success on existing public benchmarks, as Fig. 3 shown, the state-of-the-art detector, PatchCraft, suffers from significant performance drops. Moreover, the results presented in Table 4 unveil a significant performance drop across all methods. Most approaches struggle to achieve an average accuracy of around 50%, equivalent to random guessing, indicating their failure under this setting.
@@
Although our method achieves state-of-the-art results on publicly available datasets, its performance on Chameleon remains unsatisfactory. This also validates the fact that our own dataset, Chameleon, which can challenge human perception, represents a genuine problem that needs to be addressed in this field.
@@
Table 5: The classification accuracy (%) averaged over 16 test sets in $\mathcal{B}_{1}$ with specific perturbation.
@@
CNNSpot [62] 64.03 62.26 68.39 67.26
FreDect [20] 66.95 67.45 65.75 66.48
Fusing [32] 62.43 61.39 68.09 66.69
GramNet [40] 65.47 64.94 68.63 68.09
LNP [37] 53.58 54.09 67.91 66.42
LGrad [59] 51.55 51.39 71.73 69.12
DIRE-G [63] 66.49 66.12 64.00 63.09
UnivFD [49] 74.10 74.02 70.31 68.29
PatchCraft [70] 72.48 71.41 75.99 74.90
Ours 75.54 74.21 81.88 80.35

### v2

@@ 3.2.3 Dataset Annotation
At this stage, we establish an annotation platform and recruit 20 human workers to manually label each of the AI-generated images for their category and realism. For categorization, annotators are instructed to assign each image to one of four major categories: human, animal, object, and scene. Regarding realism assessment, workers are tasked with labeling the images as either Real or AI-generated, based on the criterion of “whether this image could be taken with a camera”. It’s important to note that as the annotators are not informed whether the images are generated by AI algorithms beforehand. Each image was assessed independently by two annotators, and those have been misclassified as real by both annotators can thus be considered to pass the “perception turing test” and labeled as “highly realistic”. Subsequently, we retain only those images judged as “highly realistic”. Similarly, for real images, we follow the same procedure, retaining only those belonging to the four predefined categories, as we have done for AI-generated images.
3.2.4 Dataset Comparison
Our objective is to construct a sophisticated and exhaustive test dataset for AI-generated image detection. In Table 2, we conduct a comparative analysis between our Chameleon dataset and existing datasets. Our dataset is characterized by three pivotal features: Magnitude. Comprising about 26,000 test images, the Chameleon dataset represents the most extensive collection available, enhancing its robustness. Variety. Our dataset incorporates images from a vast array of real-world scenarios, surpassing the limited categorical focus of other datasets. Resolution. With resolutions spanning from 720P to 4K, With image resolutions ranging from 720P to 4K, artifacts demand more nuanced analysis, thus presenting additional challenges to the model due to the need for fine-grained discernment. In summary, our dataset offers a more demanding and pragmatically relevant benchmark for the advancement of AI-generated image detection methodologies.
Table 2: Comparison of AI-generated image detection testset. Our Chameleon dataset is the first to encompass real-life scenarios for detector evaluation. Compared to AIGCDetectBenchmark (Zhong et al., 2023), Chameleon offers greater magnitude and superior quality, rendering it more realistic in evaluation.
ProGAN
StyleGAN
BigGAN
CycleGAN
StarGAN
GauGAN
StyleGAN2
WFIR
ADM
Glide
Midjourney
SD v1.4
SD v1.5
VQDM
Wukong
DALLE2
Chameleon
[…]
Resolution 256 256 256 256 256 256 256 1024 256 256 1024 512 512 256 512 256 720P-4K
Variety LSUN LUSN ImageNet ImageNet CelebA COCO LSUN FFHQ ImageNet ImageNet ImageNet ImageNet ImageNet ImageNet ImageNet ImageNet Real-life
4 Methodology
@@ 4.1 Patchwise Feature Extraction
To acquire the highest and lowest image patches, we use the complexity of the frequency components as an indicator. From this, we design a simple yet effective scoring mechanism. Specifically, we design $K$ different band-pass filters:
@@
where $c$ is the number of patch channels. In this way, we acquire the grades of all patches $G=\{G^{1},G^{2},...,G^{n}\}$. We then sort them to identify the highest and lowest frequency patches.
@@
Patchwise Feature Encoder. Next, firstly, these patches are resized to a size of $256\times 256$. Secondly, they are input into the SRM (Fridrich & Kodovsky, 2012) to extract their noise pattern. Lastly, these features are input into two ResNet-50 (He et al., 2016) networks ($f_{1}(\cdot)$ and $f_{2}(\cdot)$) to obtain the final feature map $F_{\text{max}}=\{f_{1}(X_{\text{max}_{1}}),f_{1}(X_{\text{max}_{2}}),...,f_{1}(X_{\text{max}_{k}})\}$, $F_{\text{min}}=\{f_{2}(X_{\text{min}_{1}}),f_{2}(X_{\text{min}_{2}}),...,f_{2}(X_{\text{min}_{k}})\}$. We acquire the highest frequency embedding and lowest frequency embedding on the mean-pooled feature:
@@ 4.2 Semantic Feature Embedding
To capture the rich semantic features within images, such as object co-occurrence and contextual relationships, we compute the visual embedding for input image with an off-the-shelf visual-language foundation model. Specifically, we adopt the ConvNeXt-based OpenCLIP model (Ilharco et al., 2021) to get the final feature map ($v\in\mathbb{R}^{h\times w\times c}$). To capture the global contexts, we append a linear projection layer followed by mean spatial pooling:
$F_{\text{s}}=\text{avgpool}(g(v)).$ (6)
@@ 4.3 Discriminator
To distinguish between AI-generated images and real images, we utilize a mixture-expert-model for the final discrimination. At low-level, we take the average of the highest frequency featured:
$F_{\text{mean}}=\text{avgpool}(F_{\text{max}},F_{\text{min}}).$ (7)
Then, we channel-wisely concatenate the representations between it and high-level embedding $F_{s}$. At last, the features are encoded into MLP to acquire the score,
$y=f([\text{avgpool}(F_{\text{mean}};F_{\text{s}}]),$ (8)
where $f(\cdot)$ denotes the MLP consisting of a linear layer, GELU (Hendrycks & Gimpel, 2016) and classifier, [;][;][ ; ] refers to the operation of channel-wise concatenation.
@@ 5 Experiments
5.1 Experimental Details
@@
Detectors. We evaluate 9 off-the-shelf detectors including CNNSpot (Wang et al., 2020), FreDect (Frank et al., 2020), Fusing (Ju et al., 2022), LNP (Liu et al., 2022a), LGrad (Tan et al., 2023), UnivFD (Ojha et al., 2023), DIRE (Wang et al., 2023), PatchCraft (Zhong et al., 2023) and NPR (Tan et al., 2024) for comparison.
Datasets. To comprehensively evaluate the generalization ability of existing approaches, we conduct experiments across two kinds of settings: Setting-I and Setting-II, which are summarized in Sec. 3.1. For the Setting-I setting, we evaluate the detectors on two general and comprehensive benchmarks of AIGCDetectBenchmark ($\mathcal{B}_{1}$) (Zhong et al., 2023) and GenImage (Zhu et al., 2024) ($\mathcal{B}_{2}$). For the Setting-II setting, we evaluate the detectors on our Chameleon ($\mathcal{B}_{3}$) benchmark. More details can be found in Appendix.
Implementation Details. AIDE includes two key modules: Patchwise Feature Extraction (PFE) and Semantic Feature Embedding (SFE). For PFE channel, we first patchify each image into patches and the patch size is set to be $N=32$ pixels. Then these patches are sorted using our DCT Scouring module with $K=6$ different band-pass filters in the frequency domain. Subsequently, we select two highest-frequency and two lowest-frequency patches using the calculated DCT scores. These selected patches are then resized to $256\times 256$ and extracted their noise pattern using SRM (Fridrich & Kodovsky, 2012). For SFE channel, we use the pre-trained OpenCLIP (Ilharco et al., 2021) to extract semantic features. We adopt data augmentations including random JPEG compression (QF $\sim\text{Uniform}(30,100)$) and random Gaussian blur ($\sigma\sim\text{Uniform}(0.1,3.0)$) to improve the robustness of detectors. Each augmentation is conducted with 10% probability. During the training phase, we use AdamW optimizer with the learning rate of $1\times 10^{-4}$ in $\mathcal{B}_{1}$ and $5\times 10^{-4}$ in $\mathcal{B}_{2}$, respectively. The batch size is set to $32$ and the model is trained on 8 NVIDIA A100 GPUs for only 5 epochs. Our method trains very quickly, only 2 hours are sufficient.
@@
Metrics. In accordance with existing AI-generated detection arpproaches (Wang et al., 2020; 2019; Zhou et al., 2018), we report both classification accuracy (Acc) and average precision (AP) in our experiments. All results are averaged over both real and AI-generated images unless otherwise specified. We primarily report Acc for evaluation and comparison in the main paper, and AP results are presented in the Appendix.
5.2 Comparison to State-of-the-art Models
@@
On Benchmark AIGCDetectBenchmark: The quantitative results in Table 3 present the classification accuracies of various methods and generators within $\mathcal{B}_{1}$. In this evaluation, all methods, except for DIRE-D, were exclusively trained on ProGAN-generated data.
@@
AIDE demonstrates a significant advancement over the current state-of-the-art (SOTA) approach, PatchCraft, achieving an average accuracy increase of 3.5%. UnivFD utilizes CLIP semantic features for detecting AI-generated images, proving effective for GAN-generated images. However, it shows pronounced performance degradation with diffusion model (DM)-generated images. This suggests that as generation quality improves, diffusion models produce images with fewer semantic artifacts, as depicted in Fig. 1 (a). Our approach, which integrates semantic, low-frequency, and high-frequency information at the feature level, enhances detection performance, yielding a 5.2% increase for GAN-based images and a 1.7% increase for DM-based images compared to the SOTA method.
Table 3: Comparison on the AIGCDetectBenchmark (Zhong et al., 2023) benchmark. Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). DIRE-D indicates this result comes from DIRE detector trained over fake images generated by ADM following its official setup (Wang et al., 2023). DIRE-G indicates this baseline is trained on the same ProGAN training data as others. The best result and the second-best result are marked in bold and underline, respectively.
Method
ProGAN
StyleGAN
BigGAN
CycleGAN
StarGAN
GauGAN
StyleGAN2
WFIR
ADM
Glide
Midjourney
SD v1.4
SD v1.5
VQDM
Wukong
DALLE2
Mean
CNNSpot 100.00 90.17 71.17 87.62 94.60 81.42 86.91 91.65 60.39 58.07 51.39 50.57 50.53 56.46 51.03 50.45 70.78
FreDect 99.36 78.02 81.97 78.77 94.62 80.57 66.19 50.75 63.42 54.13 45.87 38.79 39.21 77.80 40.30 34.70 64.03
Fusing 100.00 85.20 77.40 87.00 97.00 77.00 83.30 66.80 49.00 57.20 52.20 51.00 51.40 55.10 51.70 52.80 68.38
LNP 99.67 91.75 77.75 84.10 99.92 75.39 94.64 70.85 84.73 80.52 65.55 85.55 85.67 74.46 82.06 88.75 83.84
LGrad 99.83 91.08 85.62 86.94 99.27 78.46 85.32 55.70 67.15 66.11 65.35 63.02 63.67 72.99 59.55 65.45 75.34
UnivFD 99.81 84.93 95.08 98.33 95.75 99.47 74.96 86.90 66.87 62.46 56.13 63.66 63.49 85.31 70.93 50.75 78.43
DIRE-G 95.19 83.03 70.12 74.19 95.47 67.79 75.31 58.05 75.78 71.75 58.01 49.74 49.83 53.68 54.46 66.48 68.68
DIRE-D 52.75 51.31 49.70 49.58 46.72 51.23 51.72 53.30 98.25 92.42 89.45 91.24 91.63 91.90 90.90 92.45 71.53
PatchCraft 100.00 92.77 95.80 70.17 99.97 71.58 89.55 85.80 82.17 83.79 90.12 95.38 95.30 88.91 91.07 96.60 89.31
NPR 99.79 97.70 84.35 96.10 99.35 82.50 98.38 65.80 69.69 78.36 77.85 78.63 78.89 78.13 76.11 64.90 82.91
AIDE 99.99 99.64 83.95 98.48 99.91 73.25 98.00 94.20 93.43 95.09 77.20 93.00 92.85 95.16 93.55 96.60 92.77
Table 4: Comparison on the GenImage (Zhu et al., 2024) benchmark. Accuracy (%) of different detectors (rows) in detecting real and fake images from different generators (columns). These methods are trained on real images from ImageNet and fake images generated by SD v1.4 and evaluated over eight generators. The best result and the second-best result are marked in bold and underline, respectively.
Method Midjourney SD v1.4 SD v1.5 ADM GLIDE Wukong VQDM BigGAN Mean
ResNet-50 (He et al., 2016) 54.90 99.90 99.70 53.50 61.90 98.20 56.60 52.00 72.09
DeiT-S (Touvron et al., 2021) 55.60 99.90 99.80 49.80 58.10 98.90 56.90 53.50 71.56
Swin-T (Liu et al., 2021) 62.10 99.90 99.80 49.80 67.60 99.10 62.30 57.60 74.78
CNNSpot (Wang et al., 2020) 52.80 96.30 95.90 50.10 39.80 78.60 53.40 46.80 64.21
Spec (Zhang et al., 2019) 52.00 99.40 99.20 49.70 49.80 94.80 55.60 49.80 68.79
F3Net (Qian et al., 2020) 50.10 99.90 99.90 49.90 50.00 99.90 49.90 49.90 68.69
GramNet (Liu et al., 2020) 54.20 99.20 99.10 50.30 54.60 98.90 50.80 51.70 69.85
DIRE (Wang et al., 2023) 60.20 99.90 99.80 50.90 55.00 99.20 50.10 50.20 70.66
UnivFD (Ojha et al., 2023) 73.20 84.20 84.00 55.20 76.90 75.60 56.90 80.30 73.29
GenDet (Zhu et al., 2023) 89.60 96.10 96.10 58.00 78.40 92.80 66.50 75.00 81.56
PatchCraft (Zhong et al., 2023) 79.00 89.50 89.30 77.30 78.40 89.30 83.70 72.40 82.30
AIDE 79.38 99.74 99.76 78.54 91.82 98.65 80.26 66.89 86.88
@@
On Benchmark GenImage: In the experiments conducted on $\mathcal{B}_{2}$, all models were trained on SD v1.4 and evaluated across eight contemporary generators. Table 4 presents the results, illustrating our method’s superior performance over the current state-of-the-art, PatchCraft, with a 4.6% improvement in average accuracy. The architectural similarities between SD v1.5, Wukong, and SD v1.4, as noted by GenImage (Zhu et al., 2024), enable models to achieve near-perfect accuracy, approaching 100% on such datasets. Thus, evaluating generalization performance across other generators, such as Midjourney, ADM, and Glide, becomes essential. Our model demonstrates either the best or second-best performance in these cases, achieving an average accuracy of 86.88%.
@@
Table 5: Comparisons on the Chameleon benchmark. Accuracy (%) of different detectors (rows) in detecting real and fake images of Chameleon. For each training dataset, the first row indicates the Acc evaluated on the Chameleon testset, and the second row gives “fake image Acc / real image Acc” for detailed analysis.
@@
Training Dataset CNNSpot FreDect Fusing GramNet LNP UnivFD DIRE PatchCraft NPR AIDE
ProGAN 56.94 55.62 56.98 58.94 57.11 57.22 58.19 53.76 57.29 56.45
0.08/99.67 13.72/87.12 0.01/99.79 4.76/99.66 0.09/99.97 3.18/97.83 3.25/99.48 1.78/92.82 2.20/98.70 0.63/98.46
SD v1.4 60.11 56.86 57.07 60.95 55.63 55.62 59.71 56.32 58.13 61.10
8.86/98.63 1.37/98.57 0.00/99.96 17.65/93.50 0.57/97.01 74.97/41.09 11.86/95.67 3.07/96.35 2.43/100.00 16.82/94.38
All GenImage 60.89 57.22 57.09 59.81 58.52 60.42 57.83 55.70 57.81 63.89
9.86/99.25 0.89/99.55 0.02/99.98 8.23/98.58 7.72/96.70 85.52/41.56 2.09/99.73 1.39/96.52 1.68/100.00 22.40/95.06
On Benchmark Chameleon: As highlighted in Sec. 1, we contend that success on existing public benchmarks may not accurately reflect real-world scenarios or the advancement in AI-generated image detection, given that test sets are typically randomly sampled from generative models without ”Turing Test”. To address potential biases related to training setups—such as generator types and image categories—we evaluate the performance of existing detectors under diverse training conditions. Despite their high performance on existing benchmarks, as depicted in Fig. 3, the state-of-the-art detector, PatchCraft, experiences substantial performance declines. Additionally, Table 5 reveals significant performance decreases across all methods, with most struggling to surpass an average accuracy close to random guessing (about 50%), indicating a failure in these contexts.
@@
While our method achieves state-of-the-art results on available datasets, its performance on Chameleon remains lacking. This underscores that our dataset, Chameleon, which challenges human perception, represents a critical issue requiring attention in this field.
@@
Table 6: Robustness on JPEG Compression and Gaussian Blur of AIDE. The classification accuracy (%) averaged over 16 test sets in $\mathcal{B}_{1}$ with specific perturbation.
@@
CNNSpot 64.03 62.26 68.39 67.26
FreDect 66.95 67.45 65.75 66.48
Fusing 62.43 61.39 68.09 66.69
GramNet 65.47 64.94 68.63 68.09
LNP 53.58 54.09 67.91 66.42
LGrad 51.55 51.39 71.73 69.12
DIRE-G 66.49 66.12 64.00 63.09
UnivFD 74.10 74.02 70.31 68.29
PatchCraft 72.48 71.41 75.99 74.90
AIDE 75.54 74.21 81.88 80.35